1.Natalie has 8 socks in a drawer. 5 of the socks are black. 3 of the socks are white. Natalie takes out a sock at random, writes down its color, and puts it back into the drawer. Then Natalie takes out a second sock, at random, and writes down its color. A)What is the probability that both socks are black? B)What is the probability that the first sock is black and the second sock is white? C)What is the probability that both socks are white. D )What is the probability that the first sock is white and the second sock is black? E)What is the probability that the socks are the same color?

Answer by Boreal(15235)

a.(5/8)^2 or 25/64
-
b. (5/8)(3/8)=15/64
-
c. (3/8)^2=9/64
-
d. 15/64
-
d. the sum of a and c or 34/64. All of these may be converted to decimals and e can be reduced. Notice that the first four add to 1, as they should.